OpenStudy (babynini):

A partial sum of an arithmetic sequence is given. Find the sum. 1+5+9+...+401

OpenStudy (babynini):

a=1, d=4 Sn=(n/2)(a1+a2) I'm not sure how to plug in these numbers though!

OpenStudy (babynini):

@jim_thompson5910 :) do you have time?

jimthompson5910 (jim_thompson5910):

it should be Sn = (n/2)(a1 + an)

jimthompson5910 (jim_thompson5910):

an = a1 + d(n-1)

OpenStudy (babynini):

right right.

jimthompson5910 (jim_thompson5910):

First you need to find n

jimthompson5910 (jim_thompson5910):

an = a1 + d(n-1) 401 = 1 + 4(n-1) n = ??

OpenStudy (babynini):

101 ?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

a1 = 1 an = 401 n = 101 Sn = (n/2)*(a1 + an)

OpenStudy (babynini):

er so sn=20301

jimthompson5910 (jim_thompson5910):

I got the same

jimthompson5910 (jim_thompson5910):

wolfram agrees http://www.wolframalpha.com/input/?i=1%2B5%2B9%2B...%2B401

OpenStudy (babynini):

haha yay! the wolfram stamp of approval x) thanks so much.

jimthompson5910 (jim_thompson5910):

np